OpenStudy (mcm128):

Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.) μ = 39; σ = 15 P(50 ≤ x ≤ 70)

OpenStudy (welshfella):

you need to use the standard normal probability tables x1 = 50 standard z score z1 = 50 - 39 -------- = 0.7333 15 now calculate z2 in a similar way (x2 = 70)

OpenStudy (welshfella):

you can then read off the area under the graph from the tables using the z scores

OpenStudy (welshfella):

z score = x - m ------ s where m = mean and s = standard deviation.

OpenStudy (mcm128):

Thank you for answering! I calculated (70-39)/15 = 2.0666.

OpenStudy (mcm128):

.7333 = .62 on the Z table